Why would smaller firms be content to let a large firm practice dominant firm price leadership in an industry.

The demand equation for a product is given by:

Where I is income and P is price.

Write an equation for the point price elasticity. For what values of I and P is demand unitary elastic? Explain. (4 Marks)

Write an equation for the point income elasticity. For what values of I and P is the good a necessity? Explain. (4 Marks)

Given the total cost function for Mizoram Enterprises:



Determine the average cost function and the rate of output that will minimize average cost. (3 Marks)

Determine the marginal cost function and rate of output that will minimize marginal cost. (3 Marks)

At what rate of output does average cost equal marginal cost.


1
Expert's answer
2021-02-09T07:03:14-0500

1)The leading firm knows the industry demand curve and the supply curve of the competitive environment, it assigns its preferred price as the current market price and allows competitors to trade freely at this price.


 2)equation for the point price elasticity

"E(I)=\frac{\frac{Q2-Q1}{Q1}}{\frac{I2-I1}{I1}}"

When the coefficient of elasticity of demand for income is positive, equal to one E(I)^D=1, then the product or service is called a second-necessity product, and the demand for goods or services increases as the consumer's income increases.

When the coefficient of elasticity of demand for income is positive E(I)^D>0, the product or service is called normal, and the growth of demand for this product or service is determined by the growth of the consumer's income.


3)When marginal cost (MC) below the average total costs (ATC), the latter is reduced when MC is above ATC, the last increase

"АТС = \frac{ТС}{Q} = \frac{(FC + VC)}{Q} = (\frac{FC}{Q}) + (\frac{VC}{Q})"

"МС = \frac{\DeltaТС}{ \Delta Q}"

The MC curve also intersects the ATC curve at the point corresponding to the minimum value of the average total costs: MC=ATC